Suppose 49. g of butane is mixed with 150. g of oxygen. Calculate the minimum mass of butane that could be left over by the chemical reaction. Be sure your answer has the correct number of significant digits.
Answer #1

2C4H10 + 13 O2 ---------> 8CO2 + 10H2O  

according to balanced reaction

13 x 32 g O2 reacts with 2 x 58 g C4H10

150 g O2 reacts with 150 x 2 x 58 / 13 x 32 = 41.83 g C4H10.

mass of C4H10 left = 49 - 41.83 = 7.2 g